# Enhancing Sum Capacity via Quantum and No-Signaling Cooperation Between Transmitters

### v1

@@ III-B2 Achievability
III-C Numerical evaluation of enhancements
Although Theorem 2 suffices to show the enhancements of sum capacity (i.e., a gap is strictly positive) of a pseudo-telepathy game channel, it does not show how large the gap can be. In analyzing the gap, it suffices to evaluate $\mathcal{C}_{s}$ since $\mathcal{C}_{s}^{\mathrm{C}}=\mathcal{C}_{s}$ by Lemma 4, and Proposition 5 provides the closed-form expression of $\mathcal{C}_{s}^{\mathrm{Q}}$ or $\mathcal{C}_{s}^{\mathrm{NS}}$.
Nonetheless, computing $\mathcal{C}_{s}$ remains challenging because no single-letter characterization of the sum capacity for a general DM-IC is known. Therefore, we consider the upper bound $\overline{\mathcal{C}_{s}}$ on $\mathcal{C}_{s}$, where
$\overline{\mathcal{C}_{s}}(P_{Y_{[K]}|X_{[K]}}):=\max\limits_{P_{X_{1}}P_{X_{2}}\cdots P_{X_{K}}}I(X_{[K]};Y_{[K]}).$ (42)
Note that $\overline{\mathcal{C}_{s}}$ is the sum capacity of a DM-MAC $P_{\tilde{Y}|X_{[K]}}$ with $\tilde{Y}=Y_{[K]}$ [28, 29], and hence it provides an upper bound on the sum capacity of a DM-IC $P_{Y_{[K]}|X_{[K]}}$. Furthermore, $\overline{\mathcal{C}_{s}}$ can be numerically approximated using the generalized Blahut-Arimoto algorithm [23]. This algorithm is iterative, and its output converges from below to $\overline{\mathcal{C}_{s}}$ as the number of iterations increases, independent of the initial point. Hence, we can numerically compute the lower bound $\mathcal{C}^{r}_{s}-\overline{\mathcal{C}_{s}}$ on the gap $\mathcal{C}^{r}_{s}-\mathcal{C}_{s}$.
We apply this approach to the following examples. We consider the game channels of the form that described in Example 1 with parameters $\eta_{w}=1-\eta$ and $\eta_{l}=\eta$, for $\eta\in[0,1/2]$. The game under consideration are the CHSH, $3$-$P$, and $4$-$P$ games introduced in Section II-C. In Fig. 2, we plot the values of $\mathcal{C}_{s}^{\mathrm{Q}}$ or $\mathcal{C}_{s}^{\mathrm{NS}}$, and $\overline{\mathcal{C}_{s}}$ for the those channels. Here, we compute $\mathcal{C}_{s}^{\mathrm{Q}}$ and $\mathcal{C}_{s}^{\mathrm{NS}}$ based on the closed-form expression in Proposition 5, and $\overline{\mathcal{C}_{s}}$ using the generalized Blahut-Arimoto algorithm. The algorithm is run from $50$ randomly chosen initial points, and the maximum result is taken to reduce the effect of numerical errors. The code for this simulation is provided in the Supplementary Material.
Figure 2: Examples of the enhancements of sum capacity described in Section III-C.
For all the game channels used for the simulation, the lower bounds $\mathcal{C}^{r}_{s}-\overline{\mathcal{C}_{s}}$ on the gaps decrease as $\eta$ increases, since the game channels for $x_{[K]}\in\mathcal{W}$ become more similar to those for $x_{[K]}\not\in\mathcal{W}$ as $\eta$ increases. The values of $\mathcal{C}^{r}_{s}-\overline{\mathcal{C}_{s}}$ attain their maximum at $\eta=0$, which are approximately $0.6204$, $0.3867$, and $0.5178$ for the $4$-$P$, 33-$P$, and CHSH games, respectively.
IV Conclusion
@@
This result may provide a justification for the practical implementation of quantum-assisted communication systems. However, a limitation is that it remains unclear whether the proposed class of channels includes channels relevant to current practical communication systems. A promising direction for future work would be to determine whether quantum cooperation can enhance the sum capacity of practical channels, such as Gaussian ICs or MACs. We conjecture that the sum capacity of a Gaussian channel can be enhanced when the structure of channel gains between transmitters and receivers resembles that of a quantum pseudo-telepathy game.
Another research direction would be to derive a single-letter characterization of the capacity region of general DM-ICs or DM-MACs with cooperation between transmitters. In this context, [16] derived a multi-letter characterization of the capacity region of a DM-MAC, and [15] proposed multi-letter inner and outer bounds for the capacity region of a DM-IC. Compared to single-letter characterizations, multi-letter characterizations have the limitation that it is not clear how to compute them [24]. A single-letter characterization would help characterize the set of all channels in which cooperation enlarges the capacity region.

### v2

@@ III-B2 Achievability
IV Conclusion & Discussions
@@
Let us also address interesting future directions. First, it would be worth quantifying to what extent the sum capacity can be increased by quantum or no-signaling cooperation. For this purpose, it is necessary to quantify the sum capacity without cooperation. As one approach, the generalized Blahut-Arimoto algorithm [rezaeianComputationTotalCapacity2004], proposed for computing the classical sum capacity of DM-MACs (which serves as an upper bound on the sum capacities of DM-ICs), could be employed. However, although this algorithm can evaluate the sum capacities of certain DM-MACs, it generally provides only lower bounds on the sum capacities [buhlerNoteCapacityComputation2011]. Subsequently, [seshadriSeparationCorrelationAssistedSum2023] proposed an algorithm that computes the sum capacities of two-transmitter DM-MACs within a given additive precision in quasi-polynomial time. Consequently, it would also be interesting to investigate efficient algorithms for computing the sum capacity in multi-transmitter scenarios.
Second, it would be desirable to demonstrate sum capacity enhancements for channels relevant to current practical communication systems. A promising direction for future work would be to determine whether quantum cooperation can enhance the sum capacity of practical channels, such as Gaussian ICs or MACs. We conjecture that the sum capacity of a Gaussian channel can be enhanced when the structure of channel gains between transmitters and receivers resembles that of a quantum pseudo-telepathy game.
Another research direction would be to derive a single-letter characterization of the capacity region of general DM-ICs or DM-MACs with cooperation between transmitters. In this context, [peregMultipleAccessChannelEntangled2025] derived a multi-letter characterization of the capacity region of a DM-MAC, and [hawellekInterferenceChannelEntangled2025] proposed multi-letter inner and outer bounds for the capacity region of a DM-IC. Compared to single-letter characterizations, multi-letter characterizations have the limitation that it is not clear how to compute them [elgamalNetworkInformationTheory2011]. A single-letter characterization would help characterize the set of all channels in which cooperation enlarges the capacity region.
@@ Acknowledgments
[Uncaptioned image] Seung-Hyun Nam (Member, IEEE) received the B.S. and M.S. degrees in Electrical Engineering from Pohang University of Science and Technology (POSTECH), Pohang, South Korea, in 2018 and 2020, respectively, and Ph.D. degree in Electrical Engineering from Korea Advanced Institute of Science and Technology (KAIST), Daejeon, South Korea, in 2024. He is currently a post-doctoral researcher with Information & Electronics Research Institute, KAIST. His research interests include (both classical and quantum) information theory, statistical inference, differential privacy, and information theoretic security.
[Uncaptioned image] Hyun-Young Park (Graduate Student Member, IEEE) received the B.S. degree in Electrical Engineering from Korea Advanced Institute of Science and Technology (KAIST), Daejeon, South Korea. He is currently pursuing the Ph.D. degree in the School of Electrical Engineering at Korea Advanced Institute of Science and Technology (KAIST), Daejeon, South Korea. His research interests include information theory, differential privacy, and quantum information theory.
[Uncaptioned image] Jiyoung Yun received the B.S. and M.S. degrees in Applied Mathematics from Hanyang University, and the Ph.D. degree in Electrical Engineering from the Korea Advanced Institute of Science and Technology (KAIST). She is currently a postdoctoral researcher at KAIST. Her research interests include quantum information theory, network quantum information theory, classical and quantum channel capacities, and entanglement-assisted communication.
[Uncaptioned image] Ashutosh Rai (Member, IEEE) did his Ph.D. work [2009-2013] in theoretical physics from the S. N. Bose National Center for Basic Sciences (University of Calcutta, Kolkata, India). He has worked as a researcher at the University of Latvia (Latvia), International Institute of Physics (Brazil), and Slovak Academy of Sciences (Slovakia). He is currently working at the School of Electrical Engineering, Korea Advanced Institute of Science and Technology (KAIST). His research interest cover topics in foundation of quantum mechanics and quantum information theory.
[Uncaptioned image] Si-Hyeon Lee (Senior Member, IEEE) received the B.S. (summa cum laude) and Ph.D. degrees in electrical engineering from the Korea Advanced Institute of Science and Technology (KAIST), Daejeon, South Korea, in 2007 and 2013, respectively. She is currently an Associate Professor with the School of Electrical Engineering, KAIST. She was a Postdoctoral Fellow with the Department of Electrical and Computer Engineering, University of Toronto, Toronto, Canada, from 2014 to 2016, and an Assistant Professor with the Department of Electrical Engineering, Pohang University of Science and Technology (POSTECH), Pohang, South Korea, from 2017 to 2020. Her research interests include information theory, wireless communications, statistical inference, and machine learning. She was an IEEE Information Theory Society Distinguished Lecturer (2024-2025) and a TPC Co-Chair of IEEE Information Theory Workshop 2024. She is currently an Associate Editor for IEEE Transactions on Information Theory, an Associate Editor for IEEE Transactions on Communications, and a Guest Editor for IEEE Journal on Selected Areas in Communications (Special issue on Secure Communication, Sensing, and Computation in Future Intelligent Wireless Networks).
[Uncaptioned image] Joonwoo Bae (Member, IEEE) obtained a Ph.D. in Theoretical Physics from Universitat de Barcelona & ICFO-Institute of Photonic Sciences, Barcelona in 2007. He has worked at the Korea Institute for Advanced Study (KIAS), Centre for Quantum Technologies (CQT) in Singapore, the ICFO, Freiburg Institute for Advanced Studies (FRIAS) as a Junior Fellow, and Hanyang University. He is currently with the School of Electrical Engineering, Korea Advanced Institute of Science and Technology (KAIST). His research interests include secure quantum communication, entanglement applications, open quantum systems, quantum foundations, and their practical applications.